HOW DID PEOPLE GET THE FORMULA A=1/2(B1+B2)H FROM A=BH
Answer by stanbon(75887)

HOW DID PEOPLE GET THE FORMULA A=1/2(B1+B2)H FROM A=BH
--------------
They didn't.
A=BH is the area of a parallogram.
-------------------------------------
A = (1/2)(B1 + B2)H is the area of a trapizoid.
The formula comes from adding the area of the two
triangles formed by the diagonal of a trapizoid.
A = (1/2)B1*H + (1/2)B2*H
======================================
Cheers,
Stan H.